OpenStudy (anonymous):

what adds 4 and multiplies to -3?

OpenStudy (anonymous):

anyone know??

OpenStudy (aripotta):

you mean like, two numbers that add up to 4 and multiply to -3?

OpenStudy (anonymous):

yes

OpenStudy (aripotta):

i don't think it's possible

OpenStudy (anonymous):

thats what i thought the equation is y=-x2+4x-3

OpenStudy (anonymous):

2±√7 ?

OpenStudy (anonymous):

i got the vertex i need to find the intercepts

OpenStudy (anonymous):

Ah, if the x^2 is negative, you can multiply through by a -1 to change the signs.

OpenStudy (anonymous):

Now you can use rational numbers - now the question is, what adds to -4 and multiplies to +3?

OpenStudy (anonymous):

why is it reversed now?

OpenStudy (anonymous):

The x^2 term was negative; you need to factor that out before factoring the trinomial as a whole.

OpenStudy (anonymous):

the two isnt negative there's a -1 in front of the x^2

OpenStudy (anonymous):

That's correct. That's what I said.

OpenStudy (anonymous):

unless do i set it to -x^2+4x-3=0?

OpenStudy (anonymous):

It's the same thing. Ultimately, if you are trying to find x-intercepts, then you set the polynomial to zero, and since your leading coefficient is -1, it's easiest to multiply both sides of the equation by -1 to make it positive, then you can use the factoring shortcut.

OpenStudy (anonymous):

but then what are the two numbers? i cant even get it?

OpenStudy (anonymous):

3 is a prime number, it only has 2 factors.

OpenStudy (anonymous):

then its x-3, x-1 but i still dont know why the signs get reversed

OpenStudy (anonymous):

Here is how to do it without reversing the signs. Given -x^2+4x-3 and you want to factor it without factoring out the -1 first: You first multiply -1 by -3 to get 3. Factors of 3 that add to 4 are 1 and 3. Now split the middle term using those factors: -x^2 +x +3x -3. Now do factoring-by-grouping: x(-x+1) -3(-x+1) = (x-3)(-x+1). Still free to factor out the -1 and get -(x-3)(x-1). Personally, I'd rather not go to that trouble if I didn't have to; I'd just factor out the -1 immediately and dispose of it because it doesn't change what the x-intercepts are.